OpenStudy (anonymous):

dtermine the distance: x=7y-3 (1,-1)

OpenStudy (anonymous):

is this question complete?

OpenStudy (dumbcow):

oh the distance from the point to the line Find perpendicular line that goes through (1,-1) then find point where 2 lines intersect. use distance formula to find distance between this point and (1,-1)

OpenStudy (anonymous):

is it 11/(5sqrt(2)) ?

OpenStudy (dumbcow):

well that would give you the shortest distance,

OpenStudy (dumbcow):

i get \[d = \frac{\sqrt{11^{2}+77^{2}}}{50}\approx 1.55\]

OpenStudy (akshay_budhkar):

well the answer is 11/sqrt(50)... The formula is [x1-7y1+3]/sqrt[1+7^2]..... where x1,y1 is the point,, this is for this particular case.

OpenStudy (dumbcow):

cool i didn't know about the formula either way its the same answer :)

OpenStudy (anonymous):

x=7y-3 y= x/7 +3/7 then the normal of the line is y=(-7)x+c as, this normal goes through (1,-1) hence y=-7x+6 they meet in(.78,.54) .78-1=-.22 .54+1=1.54 hence\[\sqrt{\left( .22^{2}+1.54^{2} \right)}\]

OpenStudy (anonymous):

may b 1.55